How do u know whether the answer is negitive or postive when adding, subtracting, multiplying and dividing integers?
Found 2 solutions by Abbey, JimMarshall:
Answer by Abbey(339)

when multiplying or dividing use the odd/even rule....
Count the number of negative signs in the multiplication/division:
If there are an odd number of negative signs the answer is negative.
If there are an even number of negative signs the answer is positive.
Example:
-1 * -3 * -2 = -6 (three negatives... answer is negative)
-1 * -3 * 2 = 6 (two negatives.... answer is positive)

In addition/subtraction:
If they are both the same sign, Add the two numbers and the answer will be the same sign as both numbers
If they are different signs, subtract the smaller from the larger and the answer will have the sign of the larger number.
Example:
Same sign:
-3 - 5 = -8
3 + 5 = 8
Different Signs:
3 - 5 = -2
-3 + 5 = 2

Answer by JimMarshall(82)


when multiplying or dividing use the odd/even rule....
Count the number of negative signs in the multiplication/division:
If there are an odd number of negative signs the answer is negative.
If there are an even number of negative signs the answer is positive.
Example:
-1 * -3 * -2 = -6 (three negatives... answer is negative)
-1 * -3 * 2 = 6 (two negatives.... answer is positive)
In addition/subtraction:
If they are both the same sign, Add the two numbers and the answer will be the same sign as both numbers
If they are different signs, subtract the smaller from the larger and the answer will have the sign of the larger number.
Example:
Same sign:
-3 - 5 = -8
3 + 5 = 8
Different Signs:
3 - 5 = -2
-3 + 5 = 2